OpenStudy (anonymous):

Express the complex number in trigonometric form. -2

OpenStudy (anonymous):

MEDAL WILL BE AWARDED

OpenStudy (anonymous):

just the real number \(-2\) ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

since \(\cos(\pi)=-1, \sin(\pi)=0\) you can use \[2\left(\cos(\pi)+i\sin(\pi)\right)\]

OpenStudy (anonymous):

but the answer has to be in degrees @satellite73